# Diffusion differentiable resampling

### v1

@@ 2.3 Exponential integrators
Provided that $A$ is invertible, Jentzen2009 propose an exponential integrator
@@
where $\Delta_{k}\coloneqq t_{k}-t_{k-1}$, and the Wiener integral $B_{k}$ simplifies to $B_{k}\sim\mathrm{N}\bigl(0,\Sigma_{N}^{-1}\,(\mathrm{e}^{2\,A\,\Delta_{k}}-I_{d})\bigr)$. The integrator works effectively if the stiffness of the linear part dominates that of the non-linear part (see conditions in e.g., Buckwar2011). Indeed, the structure of the approximate score $s_{N}$ in Equation (6) is essentially a product between a Softmax function and a linear one, which is Lipschitz. However, we note the Lipschitz constant is not uniform for all $t>0$, resulting in explosive $s_{N}$ near $t=0$, such as with $V_{t}$. This exponential integrator has been empirically shown to work well for generative diffusion models in practice, for instance by Lu2025expn.
In the case when the matrix $A$ is not invertible, which rarely happens for Gaussian $\pi_{\mathrm{ref}}$ but still possibly numerically, one can also use another lower-order integrator by Lord2004:
@@
which was also considered by Zhang2023fast.
In light of Remark 1, it is even possible to simulate the resampling SDE fully in continuous time (see, e.g., Schauer2017; Baker2024), although most of the currently established techniques are still not (yet) pragmatic enough compared to just using a fine discretisation.
3 Consistency analysis
@@
Here, we have access to the weighted samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ from $\pi$ independent of the considered SDEs. We denote the distributions of $U(t)$ and $\widetilde{U}(t)$ by $q_{t}=p_{T-t}$ and $\widetilde{q}_{t}$, respectively, and similarly denote the true and approximate re-sample by $U(T)$ and $\widetilde{U}(T)$. Clearly, there are two sources of errors: 1) the ensemble score approximation $\operatorname{\nabla\!}\log p_{t}\approx s_{N}$, and 2) the initial distribution approximation $p_{T}\approx\pi_{\mathrm{ref}}$ due to finite time horizon $T$. For clarity, we here focus on continuous-time analysis, although discretisation errors may be considered within a similar framework (see e.g., Lord2014). We aim to analyse the geometric distance between the resampling distribution $\widetilde{q}_{t}$ and the target $\pi$ under these errors in relation to $t$ and $N$.
Unless otherwise needed, for any $T>0$ we assume the usual textbook linear growth and Lipschitz conditions on $X$ and $U$ so that a strong solution exists and $\int^{t}_{0}\lvert X(\tau)\rvert\mathop{}\!\mathrm{d}\tau$ has finite variance, see, for instance, Karatzas1991 or Oksendal2003. This also ensures a smooth transition density $p_{t|\tau}$ for all $0\leq\tau<t\leq T$ so that the ensemble score $s_{N}$ and its gradient in Equation (6) are pointwise well defined. We also assume the existence of the reverse process in the sense of Anderson1982, i.e., the reversal solves the same Kolmogorov forward equation in reverse time, although the established conditions for this are still implicit (see, e.g., Haussmann1986; Millet1989). We denote the earth mover’s Wasserstein distance by $\mathsf{W}_{1}(p,q)=\operatorname*{inf\,}_{\gamma\in\Gamma(p,q)}\operatorname{\mathbb{E}}[\lvert X-Y\rvert$] for $(X,Y)\sim\gamma$, where $\Gamma$ is the set of all couplings of $(p,q)$. The assumptions that we globally use are as follows.
@@
There exits constants
$\begin{split}\bigl\lvert\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(x)-\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(x^{\prime})\bigr\rvert&\leq C_{\mathrm{ref}}\,\lvert x-x^{\prime}\rvert,\\ \bigl\langle\operatorname{\nabla\!}\log p_{t}(x)-\operatorname{\nabla\!}\log p_{t}(x^{\prime}),x-x^{\prime}\bigr\rangle&\geq C_{p}\,\lvert x-x^{\prime}\rvert^{2},\end{split}$
for all $x,x^{\prime}\in\mathbb{R}^{d}$ and $t\geq 0$.
@@
There exist a constant $r>0$ and a positive non-increasing function $t\mapsto C_{e}(t)$, such that
$\operatorname*{sup\,}_{x\in\mathbb{R}^{d}}\operatorname{\mathbb{E}}\bigl[\lvert\operatorname{\nabla\!}\log p_{t}(x)-s_{N}(x,t)\rvert\bigr]\leq\frac{C_{e}(t)}{N^{r}}.$
@@
Recall that the ensemble score defined in Equation (6) is exactly a self-normalised importance sampling, where $\pi$ is the proposal, $\pi(x_{0}{\;|\;}x_{t})\propto p_{t|0}(x_{t}{\;|\;}x_{0})\,\pi(x_{0})$ is the target, and $\operatorname{\nabla\!}\log p_{t|0}(x_{t}{\;|\;}x_{0})$ is the test function. The condition in Assumption 2 is thus the same as a common non-asymptotic variance bound of importance sampling, and this has been well established by, for example, Agapiou2017 and ChopinBook2020. Typically, the Monte Carlo order is r=1/ 2r=1\,/\,2. Since the ensemble score may be unbounded as $t\to 0$, we allow $C_{e}(t)$ to depend on $t$ without further imposing a specific decay rate. Similar assumptions have been used in DeBortoli2022convergence and DeBortoli2025convergence. We obtain the following results.
@@
$\begin{split}\mathsf{W}_{1}(\widetilde{q}_{t},q_{t})&\leq\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})\,\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t}\\ &\quad+2\,b^{2}\,N^{-r}\overline{C}_{e}(t,T),\end{split}$
@@
Proposition 1 quantifies how the two errors due to score approximation and $p_{T}\approx\pi_{\mathrm{ref}}$ contribute to the resampling error. For a fixed $t$, the score error diminishes as $N\to\infty$ at the same rate $r$ but the resampling error has a bias due to $\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})$ which does not vanish. Conversely, if we fix $N$ while increasing $t$, the error bound increases exponentially in $t$, since the SDE may accumulate the score error over time. This suggests that $N$ should increase fast enough as a function of $t$ to compensate the two errors. A more specific result is obtained as follows.
@@
$\begin{split}\mathsf{W}_{1}(\widetilde{q}_{t},q_{t})&\leq 2\,b^{2}\,N^{-c}\\ &\quad+\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t-b^{2}\,C_{\mathrm{ref}}\,T}\,\mathsf{W}_{1}(\pi,\pi_{\mathrm{ref}}).\end{split}$
Furthermore, there exists a linear sequence $t\mapsto T(t)$ such that $\operatorname*{lim\,}_{t\to\infty}\mathsf{W}_{1}(\widetilde{q}_{t},q_{t})=0$.
@@
Corollary 1 shows that the diffusion resampling converges as $t\to\infty$, noting that $N$ is a function of $t$. The required assumptions are rather mild, and notably, we did not assume any specific decaying rate on $C_{e}$. The result also suggests that we should choose the reference $\pi_{\mathrm{ref}}$ close to $\pi$. Furthermore, when $t\mapsto T(t)$ is linear, the convergence (not the speed) is independent of $C_{\mathrm{ref}}$, implying that it is a plausible method, while the convergence rate largely depends on $C_{\mathrm{ref}}-2\,C_{p}$. There is likely an optimal $b$ (Lambert function) that minimises the error bound but its explicit value may be hard to know prior to running the algorithm. One may also note that the factor $N^{-c}$ seems suboptimal, as it becomes slower than the importance sampling rate $N^{-r}$. The factor can be further tightened, as explained in Appendix C.
@@
Suppose that $C_{\mathrm{ref}}\leq 2\,C_{p}$. It’s not difficult to check that Gaussian reference $\pi_{\mathrm{ref}}$, a useful setting, verifies all the assumptions. Moreover, $t\mapsto\overline{C}_{e}(t,T(t))$ can grow no faster than a polynomial in any of the two arguments, since here $t\mapsto C_{e}(t)$ decays exponentially. See also DeBortoli2025convergence.
Remark 2 show that the sample size $N$ should increase polynomially in $T$ (or equivalently $t$). This result improves over that of Corenflos2021 whose sample size scales exponentially in their entropy regularisation term $1\,/\,\varepsilon$, potentially implying cheaper computation.
A natural follow-up question is whether we can improve the error bound by eliminating the error due to $p_{T}\approx\pi_{\mathrm{ref}}$, that is, $\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})$ in Proposition 1. This may be achieved by using the technique in Corenflos2024FBS; Zhao2024rsta, where a forward-backward Gibbs chain $(p_{T|0},q_{T|0})$ is constructed to avoid explicit sampling from the marginal $p_{T}$. This essentially transforms the error $p_{T}\approx\pi_{\mathrm{ref}}$ into a statistical correlation of the chain. However, since Markov chains typically converge geometrically as well, the required number of chain steps is likely comparable to the diffusion time $T$ (Meyn2009; Andrieu2024). Furthermore, in practice we only have access to the approximate conditional $\widetilde{q}_{T|0}$. Nonetheless, we may still benefit computationally from the fact that a fixed $T$ can allow for moderate discretisation. We leave this for future investigation.
@@ 4 Experiments
In this section we empirically validate our method in both synthetic and real settings. The goal is fourfold. First, we show that the diffusion resampling, regardless of the differentiability, is a useful resampling method and is at least as good as commonly used resampling methods (Section 4.1). Second, we perform ablation experiments to test if the diffusion resampling provides a better estimate of the log marginal likelihood function (Section 4.2). Third and finally, we apply the diffusion resampling for learning neural network-parametrised SSMs using gradient-based optimisation, to show better model estimation (Sections 4.3 and 4.4). Within all these experiments, we focus on comparing to optimal transport (OT, Corenflos2021), Gumbel-Softmax (Jang2017categorical), and Soft resampling (Karkus2018). Our implementations are publicly available at https://github.com/zgbkdlm/diffres.
@@ 4.1 Gaussian mixture importance resampling
Table 1 shows that the diffusion resampling at $K=128$ gives the best SWD, followed by the baseline multinomial resampling and OT ($\varepsilon=0.3$). We also observe that the diffusion resampling performance highly depends on the choice of integrator and discretisation, which at $K=8$ is not superior than OT or multinomial. In terms of resampling variance for estimating the posterior mean, diffusion resampling is not as performant as OT but is still better than the baseline multinomial. OT seems to have stable resampling variance robust in $\varepsilon$. Gumbel-Softmax and the soft resampling methods are not comparable to diffusion or OT, except when they reduce to multinomial resampling numerically. Further results are given in in Appendix F.
@@
The time costs of diffusion resampling and OT are shown in Figure 1, and we have two important observations. The left figure shows that both methods roughly scale polynomially in the sample size $N$, where diffusion $(K=4)$ is better than OT $(\varepsilon=0.8)$. On the right side of the figure, both methods scale linearly in their parameters $K$ and $1\,/\,\varepsilon$, which aligns with the summary in Table 15, and the cross point shows that they have a similar computation at $K\approx 6\,/\,\varepsilon$ when $N=8,192$. We also observe a trend that when $N$ increases, the cross point moves left, showing that diffusion scales better than OT in $N$. although our current implementation has not yet been fully parallelised over $N$ (Algorithm 1, Line 1). Details are given in Appendix G.
@@ 4.2 Linear Gaussian SSM
We also observe that almost all the resampling methods encounter divergent parameter estimations to some extent (see more statistics in Appendix H). In particular, the divergence of Soft and Gumbel-Softmax is too significant to give a meaningful result, and these entries are thus marked as NaN. The reason is due to the quasi-Newton optimiser L-BFGS-B which heavily depends on stable and accurate gradient estimates (Xie2020). This shows that the diffusion and OT approaches are more applicable for generic optimisers, such as second-order ones. Therefore, in the next sections we will focus on comparison using first-order gradient-based methods.
@@ 4.3 Prey-predator model
where the parameters and the non-linear rate function $\lambda$ are detailed in Appendix I. We generated data at time $t\in[0,3]$ discretised by Milstein’s method with 256 steps. We model the dynamics entirely by a neural network, without assuming known the structure. After the neural network has been learnt, we make 100 predictions from the learnt model and then compare them to a reference trajectory generated by the true dynamics. All results are averaged over 20 independent runs, with $N=64$ particles used.

### v2

@@ 2.3 Exponential integrators
Provided that $A$ is invertible, Jentzen and Kloeden (2009) propose an exponential integrator
@@
where $\Delta_{k}\coloneqq t_{k}-t_{k-1}$, and the Wiener integral $B_{k}$ simplifies to $B_{k}\sim\mathrm{N}\bigl(0,\Sigma_{N}\,(\mathrm{e}^{2\,A\,\Delta_{k}}-I_{d})\bigr)$. The integrator works effectively if the stiffness of the linear part dominates that of the non-linear part (see conditions in e.g., Buckwar et al., 2011). Indeed, the structure of the approximate score $s_{N}$ in Equation (6) is essentially a product between a Softmax function and a linear one, which is Lipschitz. However, we note the Lipschitz constant is not uniform for all $t>0$, resulting in explosive $s_{N}$ near $t=0$, such as with $V_{t}$. This exponential integrator has been empirically shown to work well for generative diffusion models in practice, for instance by Lu et al. (2025).
In the case when the matrix AA is not invertible, which rarely happens for Gaussian $\pi_{\mathrm{ref}}$ but still possibly numerically, one can also use another lower-order integrator by Lord and Rougemont (2004):
@@
which was also considered by Zhang and Chen (2023).
In light of Remark 1, it is even possible to simulate the resampling SDE fully in continuous time (see, e.g., Schauer et al., 2017; Baker et al., 2024), although most of the currently established techniques are still not (yet) pragmatic enough compared to just using a fine discretisation.
3 Convergence analysis
@@
Here, we have access to the weighted samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ from $\pi$ independent of the considered SDEs. We denote the distributions of $U(t)$ and $\widetilde{U}(t)$ by $q_{t}=p_{T-t}$ and $\widetilde{q}_{t}$, respectively, and similarly denote the true and approximate re-sample by $U(T)$ and $\widetilde{U}(T)$. Clearly, there are two sources of errors: 1) the ensemble score approximation $\operatorname{\nabla\!}\log p_{t}\approx s_{N}$, and 2) the initial distribution approximation $p_{T}\approx\pi_{\mathrm{ref}}$ due to finite time horizon $T$. For clarity, we here focus on continuous-time analysis, although discretisation errors may be considered within a similar framework (see e.g., Lord et al., 2014). We aim to analyse the geometric distance between the resampling distribution $\widetilde{q}_{t}$ and the target $\pi$ under these errors in relation to $t$ and $N$.
Unless otherwise needed, for any parameter $T>0$ we assume the usual textbook linear growth and Lipschitz conditions on $X$ and $U$ so that a strong solution exists and $\int^{t}_{0}\lvert X(\tau)\rvert\mathop{}\!\mathrm{d}\tau$ has finite variance, see, for instance, Karatzas and Shreve (1991, pp. 289) or Øksendal (2007, Thm. 5.2.1). This also ensures a smooth transition density $p_{t|\tau}$ for all $0\leq\tau<t\leq T$ so that the ensemble score $s_{N}$ and its gradient in Equation (6) are pointwise well defined. We also assume the existence of the reverse process in the sense of Anderson (1982), i.e., the reversal solves the same Kolmogorov forward equation in reverse time, although the established conditions for this are still implicit (see, e.g., Haussmann and Pardoux, 1986; Millet et al., 1989). Denote the Wasserstein distance by $\mathsf{W}_{l}^{l}(p,q)=\operatorname*{inf\,}_{\gamma\in\Gamma(p,q)}\operatorname{\mathbb{E}}[\lvert X-Y\rvert^{l}]$ for $(X,Y)\sim\gamma$, where $\Gamma$ is the set of all couplings of $(p,q)$. We say a distribution $\nu$ is $z$-strongly log-concave if
$\bigl\langle\operatorname{\nabla\!}\log\nu(x)-\operatorname{\nabla\!}\log\nu(x^{\prime}),x-x^{\prime}\bigr\rangle\leq-z\,\lvert x-x^{\prime}\rvert^{2},$
and we denote it by $\nu\preceq z$. The assumptions that we globally use are as follows.
@@
There exist positive constants $2\,C_{p}<C_{\mathrm{ref}}<2\,C_{\mathrm{ref}}^{-}+2\,C_{p}$ such that
[…]
for all $x,x^{\prime}\in\mathbb{R}^{d}$, $\pi_{\mathrm{ref}}\preceq C_{\mathrm{ref}}^{-}$, and $p_{t}\preceq C_{p}$ for all $t\geq 0$.
@@
There exist a constant $r>0$ and a positive non-increasing smooth function $t\mapsto C_{e}(t)$, such that
$\operatorname*{sup\,}_{x\in\mathbb{R}^{d}}\operatorname{\mathbb{E}}\bigl[\lvert\operatorname{\nabla\!}\log p_{t}(x)-s_{N}(x,t)\rvert^{2}\bigr]^{1/2}\leq\frac{C_{e}(t)}{N^{r}}.$
@@
Recall that the ensemble score defined in Equation (6) is exactly a self-normalised importance sampling, where $\pi$ is the proposal, $\pi(x_{0}{\;|\;}x_{t})\propto p_{t|0}(x_{t}{\;|\;}x_{0})\,\pi(x_{0})$ is the target, and $\operatorname{\nabla\!}\log p_{t|0}(x_{t}{\;|\;}x_{0})$ is the test function. The condition in Assumption 2 is thus akin to non-asymptotic variance bounds of importance sampling, and this has been well established by, for example, Agapiou et al. (2017) and Chopin and Papaspiliopoulos (2020, Thm. 8.5). Typically, the Monte Carlo order is r=1/ 2r=1\,/\,2. Since the ensemble score may be unbounded as $t\to 0$, we allow $C_{e}(t)$ to depend on $t$ without further imposing a specific decay rate. Similar assumptions have been used in De Bortoli (2022, A3) and De Bortoli et al. (2025). We have the following results.
@@
$\begin{split}\mathsf{W}_{2}^{2}(\widetilde{q}_{t},q_{t})&\leq\mathsf{W}_{2}^{2}(p_{T},\pi_{\mathrm{ref}})\,\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t}\\ &\quad+2\,b^{2}\,N^{-r}\overline{C}_{e}(t,T),\end{split}$
@@
Proposition 1 quantifies how the two errors due to score approximation and $p_{T}\approx\pi_{\mathrm{ref}}$ contribute to the resampling error. For a fixed $t$, the score error diminishes as $N\to\infty$ at the same rate $r$ but the resampling error has an irreducible bias due to $\mathsf{W}_{2}(p_{T},\pi_{\mathrm{ref}})$. Conversely, if we fix $N$ while increasing $t$, the error bound increases exponentially in $t$, since the SDE may accumulate the score error over time. This suggests that $N$ should increase fast enough as a function of $t$ to compensate the two errors. A more specific result is thus obtained as follows.
@@
$\begin{split}\mathsf{W}_{2}^{2}(\widetilde{q}_{t},q_{t})&\leq 2\,b^{2}\,N^{-c}\\ &\quad+\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t-2\,b^{2}\,C_{\mathrm{ref}}^{-}\,T}\,\mathsf{W}_{2}^{2}(\pi,\pi_{\mathrm{ref}}).\end{split}$
Furthermore, there exists a linear choice $t\mapsto T(t)$ such that $\operatorname*{lim\,}_{t\to\infty}\mathsf{W}_{2}(\widetilde{q}_{t},q_{t})=0$.
@@
Corollary 1 shows that the diffusion resampling converges as $t\to\infty$, noting that $N$ is a function of $t$. The required assumptions are rather mild, and notably, we did not assume any specific decaying rate on $C_{e}$. The result also suggests that we should choose the reference $\pi_{\mathrm{ref}}$ close to $\pi$, and that the spectral gap of $\pi$ is not too large. There is likely an optimal $b$ (Lambert function) that minimises the error bound but its explicit value may be hard to know prior to running the algorithm. One may also note that the factor $N^{-c}$ seems suboptimal, as it becomes slower than the importance sampling rate $N^{-r}$. The factor can be further tightened, as explained in Appendix C.
@@
When choosing the reference $\pi_{\mathrm{ref}}$ to be a Gaussian, $t\mapsto\overline{C}_{e}(t,T(t))$ can grow no faster than a polynomial in any of the two arguments, since here $t\mapsto C_{e}(t)$ decays exponentially. See also De Bortoli et al. (2025, Remark 3).
Remark 2 show that the required sample size $N$ needs to grow only polynomially in $T$ (or equivalently $t$). This result improves upon that of Corenflos et al. (2021) whose sample size scales exponentially in their entropy regularisation term $1\,/\,\varepsilon$, and hence potentially implying cheaper computation.
A natural follow-up question is whether we can improve the error bound by eliminating the error due to $p_{T}\approx\pi_{\mathrm{ref}}$, that is, $\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})$ in Proposition 1. This may be achieved by using the technique in Corenflos et al. (2025); Zhao et al. (2025), where a forward-backward Gibbs chain $(p_{T|0},q_{T|0})$ is constructed to avoid explicit sampling from the marginal $p_{T}$. This essentially transforms the error $p_{T}\approx\pi_{\mathrm{ref}}$ into a statistical correlation of the chain. However, since Markov chains typically converge geometrically as well, the required number of chain steps is likely comparable to the diffusion time $T$ (Meyn and Tweedie, 2009; Andrieu et al., 2024). Furthermore, in practice we only have access to the approximate conditional $\widetilde{q}_{T|0}$. Nonetheless, we may still benefit computationally from the fact that a fixed $T$ can allow for moderate discretisation. We leave this for future investigation.
@@ 4 Experiments
In this section we empirically validate our method in both synthetic and real settings. The goal is fourfold. First, we show that the diffusion resampling, regardless of the differentiability, is a useful resampling method and is at least as good as commonly used resampling methods (Section 4.1). Second, we perform ablation experiments to test if the diffusion resampling provides a better estimate of the log marginal likelihood function (Section 4.2). Third and finally, we apply the diffusion resampling for learning neural network-parametrised SSMs using gradient-based optimisation in complex systems (Sections 4.3 and 4.4). Within all these experiments, we focus on comparing to optimal transport (OT, Corenflos et al., 2021), Gumbel-Softmax (Jang et al., 2017), and Soft resampling (Karkus et al., 2018). Our implementations are publicly available at https://github.com/zgbkdlm/diffres.
@@ 4.1 Gaussian mixture importance resampling
Table 1 shows that the diffusion resampling at $K=128$ gives the best SWD, followed by the baseline multinomial resampling and OT ($\varepsilon=0.3$). We also observe that the diffusion resampling performance highly depends on the choice of integrator and discretisation, which at $K=8$ is not superior than OT or multinomial. In terms of resampling variance for estimating the posterior mean, diffusion resampling is not as performant as OT but is still better than the baseline multinomial. OT seems to have stable resampling variance robust in $\varepsilon$. Gumbel-Softmax and the soft resampling methods are not comparable to diffusion or OT, except when they reduce to multinomial resampling numerically. Further results are given in Appendix F.
@@
The time costs of diffusion resampling and OT are shown in Figure 1, and we have two important observations. The left figure shows that both methods roughly scale polynomially in the sample size $N$, where diffusion $(K=4)$ is better than OT $(\varepsilon=0.8)$. On the right side of the figure, both methods scale linearly in their parameters $K$ and $1\,/\,\varepsilon$, which aligns with the summary in Table 17, and the cross point shows that they have a similar computation at $K\approx 6\,/\,\varepsilon$ when $N=8,192$. We also observe a trend that when $N$ increases, the cross point moves left, showing that diffusion scales better than OT in $N$. Details are given in Appendix G.
@@ 4.2 Linear Gaussian SSM
We also observe that almost all the resampling methods encounter divergent parameter estimations to some extent (see more statistics in Appendix H). In particular, the divergence of Soft and Gumbel-Softmax is too significant to give a meaningful result, and these entries are thus marked as NaN. The reason is due to the quasi-Newton optimiser L-BFGS-B which heavily depends on stable and accurate gradient estimates (Xie et al., 2020). This shows that the diffusion and OT approaches are more applicable for generic optimisers, such as second-order ones. Therefore, in the next sections we will focus on comparison using first-order gradient-based methods.
@@ 4.3 Prey-predator model
where the parameters and the non-linear rate function $\lambda$ are detailed in Appendix I. We generated data at time $t\in[0,3]$ discretised by Milstein’s method with 256 steps. We model the dynamics entirely by a neural network, without assuming known the dynamic structure. After the neural network has been learnt, we make 100 predictions from the learnt model and then compare them to a reference trajectory generated by the true dynamics. All results are averaged over 20 independent runs, with $N=64$ particles used.